OpenStudy (anonymous):

if the probability of seeing a moose in a day in a certain region of alaska is 0.39, what is the probability of not seeing any moose in a day there?

OpenStudy (anonymous):

Any guesses?

OpenStudy (kropot72):

The two events 'see a moose in a day' and 'not see a moose in a day' are mutually exclusive. P(A U B) = P(A) + P(B) The probability of seeing a moose in a day or not seeing a moose in a day is 1. Let the probability of not seeing a moose in a day = P(n) 1 = 0.39 + P(n) P(n) = 1 - 0.39